OpenStudy (aravindg):

The sides of a rhombus ABCD are parallel to the lines y=x+2 and y=7x+3.If the diagonal of the rhombus intersect at the point O(1,2) and the vertex A lies on y axis find possible coordinates of A

OpenStudy (aravindg):

pls help

sam (.sam.):



OpenStudy (aravindg):

can u say the properties of a rhombus ??

sam (.sam.):

AB, gradient is 1 CD, gradient is 7

OpenStudy (aravindg):

what you mean by gradient?what are the properties of a rhombus?

sam (.sam.):

Every rhombus has two diagonals connecting pairs of opposite vertices, and two pairs of parallel sides. Using congruent triangles, one can prove that the rhombus is symmetric across each of these diagonals. It follows that any rhombus has the following properties: -Opposite angles of a rhombus have equal measure. -The two diagonals of a rhombus are perpendicular; that is, a rhombus is an orthodiagonal quadrilateral. -Its diagonals bisect opposite angles.

OpenStudy (aravindg):

are diagonals equal ?? are all the four sides equal?

sam (.sam.):

four sides equal but from your equation i found that x=-1/6 and y=11/6

OpenStudy (aravindg):

wt abt diagonals?

sam (.sam.):

The point of intercection is (-1/6,11/6)

OpenStudy (aravindg):

are diagonals equal??or are they just perpendicular?

sam (.sam.):

diagonals perpendicular

OpenStudy (aravindg):

k now abt ma qn

OpenStudy (aravindg):

i think we need to find locus here...hw u got point of intersection

sam (.sam.):

y=x+2 and y=7x+3 x+2=7x+3 x=-1/6 y=7(-1/6)+3 y=11/6

OpenStudy (aravindg):

which point you talking about?

sam (.sam.):

(-1/6,11/6) could be A

OpenStudy (aravindg):